# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computation

### v1

@@ III Decoding of C642
In this work, we investigated three decoding methods for the C642. The first one is Knill’s method proposed for the ${C_{4}/C_{6}}$ scheme [44], which we refer to as hard-decision decoding. The second one is soft-decision decoding based on a posteriori probability calculation [59, 60]. We modified it for the C642 because it is not directly applicable to high-rate codes. The third one is our proposed method based on level-by-level minimum distance decoding, where we keep only minimum-distance codewords from measurement outcomes as candidates at each level from level 1 to the logical level. See Appendices DF for the details of the three methods.
@@
In the following, we assume that error correction is done by error-correcting teleportation (ECT) [44, 60, 61], where the decoding is the task to estimate the measurement outcomes of logical qubits from those of physical qubits in the logical Bell measurements (see Appendix C). Note that in ECT, we obtain all the syndrome information at once, no need to measure them one by one, unlike the surface code and qLDPC codes. We also assume that all the measurements are done in the $Z$ basis ${\{|0\rangle,|1\rangle\}}$, resulting in bit-value outcomes.
@@
We numerically evaluated the performance of the three decoding methods using a bit-flip error model (see Appendix G for details). In this work, stabilizer simulations in our numerical study were done using a python package called Stim [62]. The results are shown in Fig. 2. First, the exponents of the power-function fits for the hard-decision decoding are smaller than half the code distance, as in the case of the ${C_{4}/C_{6}}$ scheme [60], which shows its suboptimality. On the other hand, the exponents for the other two are near to or even exceed half the distance, showing their high performance. Second, the error thresholds are estimated at 1.1%, 1.5%, and 5.6%, respectively, and only our minimum distance decoding is comparable to the surface code (10.9%) [46] and a well-studied qLDPC (4%-rate hypergraph product) code (7.5%) [33]. Thus our proposed decoding method achieves the highest performance among the three methods.
@@
Figure 2: Performance of three decoders of the C642 for bit-flip errors. (a) Hard-decision decoder. (b) Symbol-MAP soft-decision decoder. (c) Level-by-level minimum distance decoder. The exponents are estimated by fitting a power function to the linear parts of the log-log plots. The thresholds indicated by arrows are defined by the intersection points of the level-3 and level-4 curves. See Appendices DF for the details of the three decoding methods and Appendix G for the simulation.
IV Fault-tolerant zero-state encoders of C642
@@
Hereafter, we consider a circuit-level noise model with error rate $p_{\mathrm{circ}}$. In this work, we assume the following noise model. Each of physical-qubit zero-state preparation and Z-basis measurement is accompanied by a bit flip with probability $p_{\mathrm{circ}}$, and each two-qubit (CNOT) gate is followed by one of 15 two-qubit Pauli errors with equal probability $p_{\mathrm{circ}}/15$ [8]. On the other hand, we assume no single-qubit-gate and memory errors for simplicity. This model is relevant for ion-trap [52, 53, 54] and neutral-atom [55, 56, 57, 58] systems.
@@
In such a circuit-level noise model, we cannot use the non-fault-tolerant encoders in Figs. 5d and 5e in Appendix B, because of uncorrectable correlated errors due to CNOT gates in a code block. Therefore we modify the zero-state encoder at each level as follows. In the following, we use various kinds of error detection and repeat encoding until no error is detected.
At level 1, we can use the well-known fault-tolerant GHZ-state preparation method with an ancilla qubit shown in Fig. 3a. We cannot apply this method directly to level 2, because of undetectable intra-block errors. To achieve fault tolerance with minimum effort at level 2, we propose the level-2 zero-state encoder shown in Fig. 3b, where level-1 error-detection gadgets are implemented by the flag-based method [63] (see Figs 12a and 12b in Appendix H). The transversal encoded CNOT and Hadamard gates can be performed fault-tolerantly by transversal physical gates, as shown in Figs. 5a and 5b in Appendix B. The level-3 zero-state encoder is the same as Fig. 3b if the levels are raised by one, but level-2 error-detection gadgets are implemented by the Steane method [64] (see Figs 12c and 12d in Appendix H). To improve the performance at level 4, we propose the level-4 zero-state encoder using two level-3 four-qubit GHZ states and error-detecting teleportation gadgets shown in Fig. 3c. As numerically shown below, these encoders well satisfy fault tolerance.
[…]
V Logical gate operations for C642
@@
Arbitrary encoded Pauli gates can easily be performed fault-tolerantly by physical Pauli gates (or the so-called Pauli frame [44, 61]) according to the definitions of the encoded Pauli operators. For the other gates necessary for universal computation (see Appendix A), we first consider the case where the same gate is performed on all the encoded/physical qubits in a code block, which we refer to as full transversal gates. If the same gate is performed on only some specified qubits in a code block, we call it partial transversal gates, which is harder to implement.
@@
From Figs. 5a and 5b in Appendix B, full transversal logical CNOT and Hadamard gates can be performed fault-tolerantly by full transversal physical CNOT and Hadamard gates (and physical SWAP gates) followed by ECT gadgets. We numerically evaluated the performance of the full transversal logical CNOT gates using the above noise model (see Appendix I for details). The results are shown in Fig. 4. The error threshold is estimated at 0.9%, which is comparable to the surface code (1.1%) [65] and a recently developed qLDPC quantum memory (0.8%) [39]. Also, the exponents of the power-function fits are close to half the code distance, showing that the logical CNOT gates (and the zero-state encoders in Fig. 3) well satisfy fault tolerance.
@@
Figure 4: Performance of full transversal logical CNOT gates for the C642 in the circuit-level error model. Each circle shows the error probability per logical CNOT gate for error rate $p_{\mathrm{circ}}$. Error bars are negligible compared with the symbol size except for the level-4 result at $p_{\mathrm{circ}}=10^{-3}$. The dotted lines are power-function fits to the five data points from the lowest, from which the exponents are estimated. See Appendix I for the details of the simulation and the main text for the definition of the error model.
@@
Next, we consider full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$ (see Appendix A for their definitions). It is known that these two gates can be performed using ancilla qubits in $|Y\rangle=(|0\rangle+i|1\rangle)/\sqrt{2}$ and $|H\rangle=\cos(\pi/8)|0\rangle+\sin(\pi/8)|0\rangle$, respectively, together with CNOT and Hadamard gates, where $|Y\rangle$ and $|H\rangle$ are the eigenstates of $Y$ and $H$ with eigenvalue 1, respectively [15, 66] (see Fig. 14 in Appendix J). Therefore the full transversal logical $S$ and ${R_{Y}(\pi/4)}$ require to fault-tolerantly prepare the level-$L$ logical $|Y\rangle_{L}^{\otimes K^{(L)}}$ and $|H\rangle_{L}^{\otimes K^{(L)}}$, which can be achieved by the standard state distillation methods only with the arbitrary-state encoder in Fig. 5e in Appendix B and full transversal encoded CNOT and Hadamard gates [15, 66]. For the transversal logical non-Clifford gates, however, we also need partial transversal logical Hadamard gates depending on the measurement outcomes (see Fig. 14b in Appendix J).
The partial transversal logical Hadamard gates can be performed using an ancilla register where some logical qubits are in $|+\rangle=H|0\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$ and the others in $|0\rangle$, which we refer to as a partial |+ket|$\rangle$| + ⟩ state. The partial |+ket|$\rangle$| + ⟩ state can be prepared fault-tolerantly (see Appendix K). From Figs. 14a and 14b in Appendix J, partial transversal logical phase and non-Clifford gates can also be performed using partial transversal logical Hadamard gates.

### v3

@@ III Decoders
In this work, we investigated three decoding methods for the many-hypercube codes. The first one is Knill’s method proposed for the ${C_{4}/C_{6}}$ scheme [45], which we refer to as hard-decision decoding. The second one is soft-decision decoding based on a posteriori probability calculation [64, 65]. The original soft-decision decoding is based on block-MAP (maximum a posteriori probability) decoding, which cannot be applied directly to high-rate codes such as the many-hypercube codes, because then we must calculate a posteriori probabilities for an exponentially large number of codewords. We thus modified it to symbol-MAP decoding applicable to high-rate codes. The third one is our proposed method based on level-by-level minimum distance decoding, where we keep only minimum-distance codewords from measurement outcomes as candidates at each level from level 1 to the logical level. See Appendices BD for the details of the three methods.
@@
In the following, we assume that error correction is done by error-correcting teleportation (ECT) [45, 65, 66]. This is quantum teleportation with logical qubits, where the classical feed-forward information is determined by decoding physical-qubit measurement outcomes, as shown in Fig. 3. Thus, the decoding is the task to estimate the measurement outcomes of logical qubits from those of physical qubits in the logical Bell measurements. Note that in ECT, we obtain all the syndrome information at once, no need to repeatedly measure them one by one, unlike the Shor method [1, 18] used for the surface code and qLDPC codes. Assuming that the two ancilla registers used for the quantum teleportation are reliable sufficiently, the errors in ECT come mainly from the decoding error. Also note that all the errors to be decoded are included in the physical-qubit measurement errors in the Bell measurement, and therefore in the case of ECT, decoders designed for independent physical-qubit errors are directly applicable to circuit-level noise models. In this work, we also assume that all the measurements are done in the $Z$ basis ${\{|0\rangle,|1\rangle\}}$, resulting in bit-value outcomes.
@@
We numerically evaluated the performance of the three decoding methods using a bit-flip error model (see Appendix E for details). In this work, stabilizer quantum-circuit simulations in our numerical study were done using a python package called Stim [67]. The decoding error probabilities estimated by the simulations are shown in Fig. 4. First, the exponents of the power-function fits for the hard-decision decoding are smaller than half the code distance, as in the case of the ${C_{4}/C_{6}}$ scheme [65], which shows its suboptimality. On the other hand, the exponents for the other two are near to or even exceed half the distance, showing their high performance. Second, the error thresholds are estimated at 1.1%, 1.5%, and 5.6%, respectively, and only our minimum distance decoding is comparable to the surface code (10.9%) [47] and a well-studied qLDPC (4%-rate hypergraph product) code (7.5%) [33]. Thus our proposed decoding method achieves the highest performance among the three methods.
Figure 3: Error-correcting teleportation. The input logical-qubit state encoded with the level-$L$ many-hypercube code, ${|\psi\rangle_{L}}$, on the left side is teleported with two logical all-zero states to the right side. The gray bold arrows show the flow of the classical information. The “D” boxes are the decoders which decode the physical-qubit measurement outcomes and provide logical-qubit ones. If a logical-qubit outcome is 1, the logical $Z$ or $X$ operator is performed on the corresponding logical qubit. The transversal logical CNOT and Hadamard gates are performed by transversal physical CNOT and Hadamard gates, as shown in Figs. 1D and 1E, respectively.
IV Fault-tolerant zero-state encoders
@@
Figure 4: Performance of three decoders of the many-hypercube codes for bit-flip errors. (A) Hard-decision decoder. (B) Symbol-MAP decoder. (C) Level-by-level minimum distance decoder. The exponents are estimated by fitting a power function to the linear parts of the log-log plots. The thresholds indicated by arrows are defined by the intersection points of the level-3 and level-4 curves. See Appendices BD for the details of the three decoding methods and Appendix E for the simulation.
Figure 5: Fault-tolerant zero-state encoders of the many-hypercube codes. (A) Level-1 zero-state encoder. The inputs are seven physical qubits in zero states. The CNOT and Hadamard gates are physical ones. The encoding is repeated until the measurement outcome of the ancilla qubit is 0. (B) Level-2 zero-state encoder. The transversal level-1 encoded CNOT and Hadamard gates are performed by transversal physical gates, as in Figs. 1D and 1E. The measurement outcomes of the level-1 encoded ancilla qubits are obtained by decoding the physical-qubit outcomes with the hard-decision decoding for error detection (see Appendix B). The encoding is repeated until no error is detected and all the level-1 outcomes are 0. (C and D) Level-1 $Z$-error and $X$-error detection gadgets (EDZL1 and EDXL1), respectively, used for the level-2 zero-state encoder in (B). Errors are detected unless both the measurement outcomes of the two ancilla qubits are 0. (E and F) Level-2 $Z$-error and $X$-error detection gadgets (EDZL2 and EDXL2), respectively, used for the level-3 zero-state encoder. Errors are detected unless none of the outcomes of the minimum distance decoding for error detection with $L_{D}=1$ is $F$ (see Appendix D). Level-3 zero-state encoder is similar to (B) except replacing EDZL1 and EDXL1 with EDZL2 and EDXL2. (G) Level-4 zero-state encoder. EDTL3 denotes a level-3 error-detecting teleportation gadget, where the minimum distance decoding for error detection with $L_{D}=2$ is used (see Appendix D). The same decoding is used for decoding the measurement outcomes of the two ancilla registers. The encoding is repeated until no error is detected and all the parities of the outcomes of the level-3 encoded-qubit pairs in the two ancilla registers are even, where if both the outcomes of a pair are 1, we perform encoded $X$ gates on either of the first or second half of the six level-3 encoded qubits.
@@
Hereafter, we consider a circuit-level noise model with error rate $p_{\mathrm{circ}}$. In this work, we assume the following noise model. Each of physical-qubit zero-state preparation and $Z$-basis measurement is accompanied by a bit flip with probability $p_{\mathrm{circ}}$, and each two-qubit (CNOT) gate is followed by one of 15 two-qubit Pauli errors with equal probability $p_{\mathrm{circ}}/15$ [8]. On the other hand, we assume no single-qubit-gate and memory errors for simplicity. This model is relevant for ion-trap [53, 54, 55, 56, 57, 58, 59] and neutral-atom [60, 61, 62, 63] systems. (Even for these systems, it is eventually desirable to consider the effect of the neglected errors, but it requires thorough optimization of time overheads and therefore is left for future work.)
@@
In such a circuit-level noise model, we cannot use the non-fault-tolerant encoders in Figs. 1A and 1B, because of undetectable correlated errors due to CNOT gates in a code block. Therefore we modify the zero-state encoder at each level as follows. In the following, we use various kinds of error detection and repeat encoding until no error is detected.
For the encoding of the level-1 all-zero state, which is the six-qubit GHZ state as shown in Eq. (5), we can use the well-known fault-tolerant GHZ-state preparation method with an ancilla qubit shown in Fig. 5A. We accept the encoding if the measurement outcome of the ancilla qubit is 0, otherwise reject and restart it from the beginning. Thus logical $X$ errors can be eliminated. (Logical $Z$ errors do not occur on the logical all-zero state.) To evaluate the space and time overheads for the zero-state encoders, here we introduce the total number of physical qubits including ancilla ones, $N^{\prime(L)}$, and the circuit depth including state preparation and measurement, $T^{(L)}$, respectively, required for the level-$L$ zero-state encoder. From Fig. 5A, we obtain $N^{\prime(1)}=7$ and $T^{(1)}=8$.
We cannot apply this method directly to level 2, because of uncorrectable intra-block errors. Also, the low-depth circuit for the GHZ-state preparation in Fig 5A results in many error-detection gadgets to eliminate the intra-block $Z$ errors. To achieve fault tolerance with minimum effort at level 2, we propose the level-2 zero-state encoder shown in Fig. 5B, where the control qubits of the CNOT gates concentrate on a single qubit in order to eliminate the intra-block $Z$ errors by a single error-detection gadget. To eliminate intra-block and logical $X$ errors, we also need only a single error-detection gadget and ancilla-qubit measurement. To minimize space overheads, the level-1 error-detection gadgets are implemented by the flag-based method with two physical ancilla qubits and depth 12 in Figs. 5C and 5[68]. The transversal encoded CNOT and Hadamard gates can be performed fault-tolerantly by transversal physical gates with depth 1 and 2, respectively, as shown in Figs. 1D and 1E. Optimizing the overlaps of physical operations to minimize the time overhead, we obtain $N^{\prime(2)}=N^{\prime(1)}\times 7+2\times 2=53$ and $T^{(2)}=T^{(1)}+2+6+(12-2)=26$. Note that the space overhead, i.e., the total number of physical qubits, can be reduced by reusing the ancilla qubits, but then the time overhead will increase. Also note that we can use the Steane method with an encoded ancilla qubit [69] for the error-detection gadgets, which in the level-2 case are shown in Figs. 5E and 5F, but this results in larger space overheads and no performance improvement (see Appendix G). This is because at level 1, we only need to measure just a single weight-6 stabilizer, which can be achieved most easily by the flag-based method.
The level-3 zero-state encoder is the same as the level-2 one in Fig. 5B if the levels are raised by one, but level-2 error-detection gadgets are implemented by the above-mentioned Steane method [69] shown in Figs. 5E and 5F. This is because at level 2, we need to measure six weight-6 level-1 stabilizers and four weight-12 level-2 stabilizers, which can be achieved by the Steane method more easily than the Shor method with physical ancilla qubits. The space and time overheads at level 3 are estimated as $N^{\prime(3)}=N^{\prime(2)}\times 7+N^{\prime(2)}\times 2=477$ and $T^{(3)}=T^{(2)}+2+6+4=38$. where and in the following, decoding is not counted for time-overhead evaluation.
Although the encoder in Fig. 5B can also be applied to the level-4 case, then the acceptance probabilities at error-detection gadgets become too low if we use the most stringent condition for error detection. To mitigate this issue, we can relax the error-detection condition (see Appendix D), but then the logical gate performance becomes low (see Appendix H for detailed results using the encoder in Fig. 5B at level 4). To improve the performance, we propose the level-4 zero-state encoder shown in Fig. 5G, which uses two level-3 four-qubit GHZ states with error-detecting teleportation (EDT) gadgets eliminating intra-block $Z$ and $X$ errors simultaneously. The reason why this encoder can achieve higher performance than that in Fig. 5B is that the intra-block errors in two four-qubit GHZ states are independent and therefore detectable with high probability by the ancilla measurements. From Fig. 5G, the space and time overheads at level 4 are estimated as $N^{\prime(4)}=N^{\prime(3)}\times 8+N^{\prime(3)}\times 2\times 2=5724$ and $T^{(4)}=T^{(3)}+2+3+4=47$. where and in the following, Pauli operations are not counted for time-overhead evaluation.
As numerically shown below, the encoders in Fig. 5 well satisfy fault tolerance. The time overheads evaluated above are almost proportional to the level. (The effect of postselection at error-detection gadgets is discussed later.) On the other hand, the logical-gate error rates decrease doubly exponentially with respect to the level, as numerically shown later. These facts suggest that the encoding only needs doubly logarithmic time overheads with respect to computational size and hence is time efficient [44]. Although the space overheads increase more rapidly, the net encoding rates defined by $K^{(L)}/N^{\prime(L)}$ including ancilla qubits are still relatively high, higher than 4% even at level 4. (The encoding rate for the same-distance surface code is $0.41/16^{2}=$ 0.4% even without ancilla qubits.) Further optimization of the space and time overheads may be possible but is left for future work.
The comparison with the $C_{4}/C_{6}$ scheme [45, 65] is as follows. The parameters of the $C_{4}/C_{6}$ scheme are given by $N^{(L)}=4\times 3^{L-1}$, $K^{(L)}=1$ (or 2), $D^{(L)}=2^{L}$, and $N^{\prime(L)}=4\times 12^{L-1}$ [65]. By the present technique for the many-hypercube codes, the space overheads may be reduced to $N^{\prime(L)}=5^{L}$. Then the net encoding rate at level 4 is $K^{(4)}/N^{\prime(4)}=0.16$% (or 0.32%), which is much lower than that of the level-4 many-hypercube code, as expected. On the other hand, the logical CNOT performance of the $C_{4}/C_{6}$ scheme [65] is much higher than that of the many-hypercube codes presented below. This may be due to the high rates of the many-hypercube codes and the optimal (block-MAP) decoder for the $C_{4}/C_{6}$ scheme.
V Logical gate operations
Arbitrary encoded Pauli gates can easily be performed fault-tolerantly by physical Pauli gates (or the so-called Pauli frame [45, 66]) according to the definitions of the encoded Pauli operators, e.g., in Eqs. (12) and (13).
Logical SWAP gates can also be performed easily by physical SWAP gates or renumbering of physical qubits as follows. The level-1 encoded SWAP gates are performed as shown in Fig. 1C. At higher levels, e.g., at level 3, simultaneous physical SWAP gates $\{Q_{1,j,k}\leftrightarrow Q_{3,j,k}|j,k=1,\ldots,6\}$ result in simultaneous logical SWAP gates $\{Q_{1,j^{\prime},k^{\prime}}^{(3)}\leftrightarrow Q_{2,j^{\prime},k^{\prime}}^{(3)}|j^{\prime},k^{\prime}=1,\ldots,4\}$. Similarly, $\{Q_{i^{\prime},3,k^{\prime}}^{(3)}\leftrightarrow Q_{i^{\prime},4,k^{\prime}}^{(3)}|i^{\prime},k^{\prime}=1,\ldots,4\}$ and $\{(Q_{i^{\prime},j^{\prime},1}^{(3)},Q_{i^{\prime},j^{\prime},2}^{(3)})\leftrightarrow(Q_{i^{\prime},j^{\prime},3}^{(3)},Q_{i^{\prime},j^{\prime},4}^{(3)})|i^{\prime},j^{\prime}=1,\ldots,4\}$ are performed by corresponding simultaneous physical SWAP gates, $\{Q_{i,4,k}\leftrightarrow Q_{i,6,k}|i,k=1,\ldots,6\}$ and $\{(Q_{i,j,1},Q_{i,j,2},Q_{i,j,3})\leftrightarrow(Q_{i,j,4},Q_{i,j,5},Q_{i,j,6})|i,j=1,\ldots,6\}$, respectively. In general, we can perform simultaneous logical SWAP gates between the logical qubits corresponding to the hypercubes on two parallel hyperplanes by simultaneous physical SWAP gates.
@@ V Logical gate operations
For the other gates necessary for universal computation (see Appendix A), we first consider the case where the same gate is performed on all the encoded/physical qubits in a code block, which we refer to as full transversal gates. If the same gate is performed on only some specified qubits in a code block, we call it partial transversal gates, which is harder to implement.
@@
From Figs. 1D and 1E, full transversal logical CNOT and Hadamard gates can be performed fault-tolerantly by full transversal physical CNOT and Hadamard gates (and physical SWAP gates) followed by ECT gadgets. Their time overheads are dominated by the ECT gadgets, which are estimated at $T^{(L)}+7$ in the level-L case from Fig. 3, and therefore time efficient. We numerically evaluated the performance of the full transversal logical CNOT gates using the above-mentioned circuit-level noise model (see Appendix F for details). The logical CNOT error probabilities estimated by the simulation are shown in Fig. 6A. The error threshold is estimated at 0.9%, which is comparable to the surface code (1.1%) [70] and a recently developed qLDPC quantum memory (0.7%) [40]. Also, the exponents of the power-function fits are close to half the code distance, showing that the logical CNOT gates (and the zero-state encoders in Fig. 5) well satisfy fault tolerance. (The level-4 value at $p_{\mathrm{circ}}=10^{-3}$ seems an outlier, resulting in a smaller exponent, e.g., 7, in the range of $p_{\mathrm{circ}}\leq 10^{-3}$ than that estimated by fitting, 7.8, shown in Fig. 6A. The exponent smaller than half the code distance may be due to the relaxation of the error-detection condition mentioned above.)
In the above simulation of logical CNOT gates, we also estimated the expectation value of the total number of physical qubits to prepare a logical all-zero state $|0\rangle_{L}^{\otimes K^{(L)}}$ taking the effect of postselection and restart into account. The results are shown in Fig. 6B. It turns out that we need $p_{\mathrm{circ}}\leq 10^{-3}$ to achieve the increasing rate less than 2 even at level 4. This will be achievable for ion-trap and neutral-atom systems [38]. [In Fig. 6B, the numbers at $p_{\mathrm{circ}}=0$ are $N^{\prime(L)}$ in Eqs. (20)–(23).]
@@
Figure 6: Performance of full transversal logical CNOT gates for the many-hypercube codes in the circuit-level error model. (A) Each circle shows the error probability per logical CNOT gate for error rate $p_{\mathrm{circ}}$. The dotted lines are power-function fits to the five data points from the lowest, from which the exponents are estimated. (B) The number of physical qubits (space overhead) required for the preparation of a logical all-zero state. See Appendix E for the details of the simulation and the main text for the definition of the error model.
@@
Next, we consider full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$ (see Appendix A for their definitions). It is known that these two gates can be performed using ancilla qubits in $|Y\rangle=(|0\rangle+i|1\rangle)/\sqrt{2}$ and $|H\rangle=\cos(\pi/8)|0\rangle+\sin(\pi/8)|0\rangle$, respectively, together with CNOT and Hadamard gates (see Figs. 7A and 7B), where $|Y\rangle$ and $|H\rangle$ are the eigenstates of $Y$ and $H$ with eigenvalue 1, respectively [15, 71]. Therefore the full transversal logical $S$ and ${R_{Y}(\pi/4)}$ require to fault-tolerantly prepare the level-$L$ logical $|Y\rangle_{L}^{\otimes K^{(L)}}$ and $|H\rangle_{L}^{\otimes K^{(L)}}$, which can be achieved by non-fault-tolerant encoding with the arbitrary-state encoder in Fig. 1B followed by state distillation and level-raising teleportation, as originally proposed for the $C_{4}/C_{6}$ scheme [71]. (The full transversal logical non-Clifford gates also need partial transversal logical Hadamard gates depending on the measurement outcomes, which are explained later.) We propose the 2-to-1 distillation method for $|Y\rangle_{L}^{\otimes K^{(L)}}$ based on $HS|Y\rangle=|1\rangle$ shown in Fig. 7C. Note that our method needs only two noisy $|Y\rangle_{L}^{\otimes K^{(L)}}$, unlike the well-known 7-to-1 distillation method with the Steane code [15]. This difference comes from the fact that the many-hypercube codes are based on distance-2 codes, and therefore do not need distance-3 codes, such as the Steane code, for distillation. For $|H\rangle_{L}^{\otimes K^{(L)}}$, we can use the standard 14-to-2 distillation method based on the distance-2 $H_{6}$ code [71, 72]. The level-raising teleportation shown in Fig. 7D can convert four distilled $|Y\rangle_{L-1}^{\otimes K^{(L-1)}}$ and $|H\rangle_{L-1}^{\otimes K^{(L-1)}}$ to $|Y\rangle_{L}^{\otimes K^{(L)}}$ and $|H\rangle_{L}^{\otimes K^{(L)}}$, respectively, in a fault-tolerant manner [71].
In the following, we consider partial transversal logical gates. The partial transversal logical Hadamard gates can be performed using an ancilla register where some logical qubits are in $|+\rangle=H|0\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$ and the others in $|0\rangle$, which we refer to as a partial |+ket|$\rangle$| + ⟩ state, together with full transversal logical CNOT and Hadamard gates and an ECT gadget, as shown in Fig. 8A. (The ECT gadget is necessary to prevent intra-block errors from spreading through the logical CNOT and Hadamard gates.) Logical partial |+ket|$\rangle$| + ⟩ states can be prepared fault-tolerantly in a similar manner to $|Y\rangle_{L}^{\otimes K^{(L)}}$ and $|H\rangle_{L}^{\otimes K^{(L)}}$ explained above. At level 1, we can fault-tolerantly prepare encoded partial |+ket|$\rangle$| + ⟩ states more efficiently than state distillation, as follows. Any level-1 encoded partial |+ket|$\rangle$| + ⟩ state can be prepared by one of the three methods in Figs. 8B–8D. The first and second encoders in Figs. 8B and 8C are based on the fact that some level-1 partial |+ket|$\rangle$| + ⟩ states can be expressed with only the Bell state, $(|00\rangle+|11\rangle)/\sqrt{2}$, and the four-qubit GHZ state by definition. The third encoder in Fig. 8D is obtained from the arbitrary encoder in Fig. 1B followed by verification with minimum effort through the measurements of an encoded $Z$ operator and an encoded $X$ operator, in a similar manner to the most efficient fault-tolerant Steane-code encoder with a single ancilla qubit [73]. Applying the level-raising teleportation to the resultant level-1 partial |+ket|$\rangle$| + ⟩ states, we obtain level-2 partial |+ket|$\rangle$| + ⟩ states. At higher levels, we use the 4-to-1 distillation method shown in Fig. 8E.
The total number of physical qubits and circuit depth for the preparation of a level-$L$ logical partial |+ket|$\rangle$| + ⟩ state, which are denoted by $N_{+}^{(L)}$ and $T_{+}^{(L)}$, respectively, are estimated from Figs. 7D, 8D, and 8E in the worst-case scenario where all the level-1 blocks are encoded by the encoder in Fig. 8D as follows: $N_{+}^{(1)}=8$, $T_{+}^{(1)}=9$, $N_{+}^{(2)}=N_{+}^{(1)}\times 4+N^{\prime(2)}\times 2=138$, $T_{+}^{(2)}=T_{+}^{(1)}+6+4=20$, $N_{+}^{(3)}=N_{+}^{(2)}\times 4\times 4+N^{\prime(3)}\times 2=3162$, $T_{+}^{(3)}=T_{+}^{(2)}+6+4=30$, $N_{+}^{(4)}=N_{+}^{(3)}\times 4\times 4+N^{\prime(4)}\times 2=62040$, and $T_{+}^{(4)}=T_{+}^{(3)}+6+4=40$. Although this preparation is time efficient, the space overheads are considerably larger than those for the encoded all-zero states. More efficient encoding of partial |+ket|$\rangle$| + ⟩ states may be possible but left for future work.
From Figs. 7A and 7B, we can perform partial transversal logical $S$ and $R_{Y}(\pi/4)$ by replacing the full transversal logical Hadamard gates with partial ones explained above, because then the two logical CNOT gates on the qubits without logical Hadamard gates are cancelled out.
Figure 7: Logical phase and non-Clifford gates. (A and B) Full transversal logical phase and non-Clifford gates, $S$ and ${R_{Y}(\pi/4)}$, respectively. In (B), the measurement outcomes are decoded, and the logical $Z$ and $H$ are performed if the corresponding decoding outcome is 1. ECT is inserted between two CNOT gates for preventing intra-block errors from spreading due to physical SWAP gates in the logical Hadamard gate. Partial transversal logical phase and non-Clifford gates can be performed by replacing the full transversal logical Hadamard gates with partial ones in Fig. 8A. (C) Proposed 2-to-1 distillation for $|Y\rangle_{L}^{\otimes K^{(L)}}$. The distillation is accepted if all the decoding outcomes are 1. (D) Level-raising teleportation. The box “Encode-1” is the gadget performing the inversion of the arbitrary-state encoder in Fig. 1B, outputting four level-($L$-1) states, where the other two level-($L$-1) blocks are measured and decoded by an error-detection decoder, and the Bell-state preparation is repeated until all the decoding outcomes are 0. In the Bell measurements, the measurement outcomes are also decoded by an error-detection decoder [71].
Figure 8: Partial transversal logical Hadamard gates. (A) Partial transversal logical Hadamard gates with a logical partial |+ket|$\rangle$| + ⟩ state and full transversal logical CNOT and Hadamard gates. The logical $Z$ ($X$) is performed if the decoding outcome is 0 (1). ECT is inserted between two CNOT gates for preventing errors from spreading due to physical SWAP gates in the logical Hadamard gate. (B–D) Fault-tolerant preparation methods for level-1 logical partial |+ket|$\rangle$| + ⟩ states. The other ones can also be prepared similarly to one of the three through SWAP gates and full transversal Hadamard gates. (E) 4-to-1 distillation for a logical partial |+ket|$\rangle$| + ⟩ state. This is repeated until all the decoding outcomes are consistent with the input state.
Figure 9: Logical CNOT gates. (A) Partial transversal logical CNOT gates (inter-block logical CNOT gates) with two ancilla registers in the same partial |+ket|$\rangle$| + ⟩ states (showing only two logical qubits) and full transversal logical CNOT and Hadamard gates. The $Z$ and $X$ are performed if the corresponding decoding outcome is 1. Then, the logical CNOT gates denoted by $U_{\mathrm{CNOT}}$ are performed on the logical qubits corresponding to |+ket|$\rangle$| + ⟩ of the partial |+ket|$\rangle$| + ⟩ state in the control and target states, $|\psi_{c}\rangle_{L}$ and $|\psi_{t}\rangle_{L}$, respectively. (B) Intra-block logical CNOT gates using an ancilla register in the logical all-zero state. The first and third CNOT gates highlighted in red are partial transversal logical CNOT gates on the control qubits of the logical CNOT gates to be performed on $|\psi\rangle_{L}$. (The second one is full transversal logical CNOT gates.) The SWAP gate moves the logical qubits corresponding to the control qubits to the ones corresponding to the target qubits of the logical CNOT gates to be performed. They are returned to the original positions by the SWAP-1 gate. Thus the desired intra-block logical CNOT gates denoted by $U_{\mathrm{CNOT}}$ are performed on $|\psi\rangle_{L}$.